OpenStudy (anonymous):

Using the substitution method and i need to show my work y=x+3 y=5x-5

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow y = x + 3 }\) We can take y and replace it with x + 3 \(\Large \color{purple}{\rightarrow x + 3 = 5x - 5 }\) Can you now solve the linear equation?

OpenStudy (callisto):

y=x+3 -(1) y=5x-5 -(2) sub. (1) into (2) x+3 = 5x-5 4x = 8 x =? y = ? + 3 = ???

OpenStudy (anonymous):

thank you for the help. i get it x=2, y=5

OpenStudy (callisto):

Yup :)